OpenStudy (anonymous):

What is the area of a circle with circumference of 9.42 inches?

OpenStudy (anonymous):

you would have to know the following formulae: \[\Large C=2\pi r\] \[\Large A=\pi r^2\] C=Circumference of the circle A=Ara of the circle r=radius of the circle

OpenStudy (anonymous):

using the first formula containing the circumference of the circle, you can be able to find the radius... after finding the radius, you could place that into the second formula to find the area of thee circle

OpenStudy (anonymous):

can you do that??